Rose's garden is in the shape of a trapezoid. If the height of the trapezoid is 16 m, one base is 20 m, and the area is 224 m^2, find the length of the other base.
Answer by favegirl13(22)

Area of trapezoid:
A=(b1+b2)*h/2
224=(20+x)*16/2 (x represents second base number which we are trying to find out)
224=20x*16/2
224=3584x/2
224=1792
224/224=1792/224
x=8
the second base # is 8
CONFIRM ANSWER:
A=(b1+b2)*h/2
224=(20+8)*16/2
224=28*16/2
224=448/2
the area is 224
so that confirms that the second base number is 8